OpenStudy (anonymous):

Integrate [(x^2)+1]/[sqrt(x)]

OpenStudy (anonymous):

\[\int\frac{x^2+1}{\sqrt x}dx=\int\frac{x^2+1}{x^{\frac{1}{2}}}dx=\int\left(x^{\frac{3}{2}}+x^{-\frac{1}{2}}\right)\;dx\]